OpenStudy (anonymous):

The midpoint of a line segment with end points as (2, 3) and (4, y2) is (3, 2). What is the value of y2?

myininaya (myininaya):

\[(\frac{2+4}{2}, \frac{3+y^2}{2})=(3,2)\]

myininaya (myininaya):

\[=>\frac{2+4}{2}=3 \text{ and } \frac{3+y^2}{2}=2\]

myininaya (myininaya):

the first part is a true statement

myininaya (myininaya):

so all you need to look at is the second part

myininaya (myininaya):

\[\frac{3+y^2}{2}=2\]

myininaya (myininaya):

multiply 2 on both sides then subtract 3 and then you will get...what?